18.02A Second Half, Spring Semester 2004 General Information

Class meetings: Tuesday and Thursday 1:00–2:00, Friday 2:00–3:00 in 4-145. Text: Simmons, Calculus with Analytic Geometry, 2nd edition, McGraw-Hill. 18.02A Supplementary Notes (CopyTech, 11-004). The notes are what you got halfway through the fall semester—no purchase necessary. Lecturer: David Vogan, 2-281 (x3-4991, dav@math.mit.edu). Office hours are Tuesday 11–12 and Wednesday 3–4, or by appointment. Recitation meetings: Tuesday and Thursday 12:00–1:00 in 4-145. Recitation instructor: S. Francisco, 2-229 (x3-1589, francisco@mit.edu). Tutoring is available in 2-102 Monday through Thursday 3:00–5:00 and 7:30–9:30. Homework assigned most Tuesdays in lecture. Due in recitation the following Tuesday. Late work will not be accepted. Exams: One-hour exams during the lecture hour on Thursday, February 26, and Thursday, March 18. The exams will be closed book, and calculators will not be allowed. Grading: The first exam will be worth 30%, the second 40%, and the problem sets 30%.

Syllabus

In the reading assignments, “V1” means section V1 of the Supplementary Notes; “21.1” means section 21.1 of Simmons. Doing the reading before class offers you the priceless (\begin{matrix}c \\ ◯ \end{matrix}MasterCard) possibility of catching the lecturer in an error.

Tues 2/3 Lecture 40 vector fields and line integrals in the plane (V1, 21.1)
Thurs 2/5 Lecture 41  gradient and conservative fields (21.2)
Fri 2/6 Lecture 42 potential functions (V2)
Tues 2/10 Lecture 43  Green’s theorem (21.3)
Thurs 2/12 Lecture 44  flux and the normal form of Green’s theorem (V3, V4)
Fri 2/13 Lecture 45 extensions and applications of Green’s theorem (V5, V6 pp. 1–2)
Tues 2/17 NO MEETING Monday classes held
Thurs 2/19 Lecture 46  triple integrals; rectangular and cylindrical coordinates (I3, 20.5, 20.6)
Fri 2/20 Lecture 47 spherical coordinates; gravitational attraction (20.7, G)
Tues 2/24 Lecture 48  applications
Thurs 2/26 Lecture 49  Exam covering 40–48
Fri 9/27  Lecture 50  vector fields in space; flux; surface integrals (V8, V9)

Tues 3/2  Lecture 51  surface integrals continued; divergence theorem (V10)
Thurs 3/4 Lecture 52  applications; ∇ notation (V15)
Fri 3/5  Lecture 53  line integrals in space; curl F and conservative fields (V11, V12)
Tues 3/9  Lecture 54  potential functions in space; Stokes’ theorem (V13)
Thurs 3/11 Lecture 55  Stokes’ theorem continued; topological considerations (V14)
Fri 3/12  Lecture 56  applications (V15)
Tues 3/16 Lecture 57  review
Thurs 3/18 Lecture 58  Exam covering 40–57